Zeta Enterprises’ common stock dividend is expected to grow at a long-run rate of 3% per year. The dividend recently paid was $1.50 per share. Investors require a 13% return from Zeta’s common stock. What is your estimate of Zeta Enterprises’ common stock price?
Answer #1

Div Po r-g Po= Price of Stock Div = Estimat ed Dividends for Next Period rRequired Rate of Return g Growth Rate

Div1 = Div0 * (1 + g)

Div1 = $1.50 * (1 + 3%) = $1.545

1.545 Ро 0.13 0.03

1.545 Ро 0.10

--> Answer Po 15.45